OpenStudy (anonymous):

The gasoline gauge on a van initially read 1/8 full. When 6 gallons were added to the tank, the gauge read 1/2 full. How many more gallons are needed to fill the tank.

OpenStudy (anonymous):

first start out by setting common denominators of 8. 1/8 and 4/8.

OpenStudy (anonymous):

@genius12 how did you get 4/8x?

OpenStudy (anonymous):

Let's call the total capacity of the tank x. Then: 1/8x + 6 = 1/2x 6 = 1/2x - 1/8x 6 = 4/8x - 1/8x 6 = 3/8x x = 6 / (3/8) x = 6 * 8/3 x = 48/3 x = 16 Thus, the tank has a capacity of 16 gallons. If it's half full after adding 6 gallons, then it's 16/2 = 8 gallons full. Since the total capacity is 16, then the tank needs 16 - 8 = 8 gallons more to be completely full.

OpenStudy (anonymous):

To subtract fractions, you make the denominators the same. So to have a denominator of 8, I multiplied the top and bottom of 1/2 by 4, which gives 4/8. @vanessa32

OpenStudy (anonymous):

@genius12 I understand, thank you